### v1

# Recognition through Reasoning: Reinforcing Image Geo-localization with Large Vision-Language Models

Abstract
Previous methods for image geo-localization have typically treated the task as either classification or retrieval, often relying on black-box decisions that lack interpretability. The rise of large vision-language models (LVLMs) has enabled a rethinking of geo-localization as a reasoning-driven task grounded in visual cues. However, two major challenges persist. On the data side, existing reasoning-focused datasets are primarily based on street-view imagery, offering limited scene diversity and constrained viewpoints. On the modeling side, current approaches predominantly rely on supervised fine-tuning, which yields only marginal improvements in reasoning capabilities. To address these challenges, we propose a novel pipeline that constructs a reasoning-oriented geo-localization dataset, MP16-Reason, using diverse social media images. We introduce GLOBE, Group-relative policy optimization for Locatability assessment and Optimized visual-clue reasoning, yielding Bi-objective geo-Enhancement for the VLM in recognition and reasoning. GLOBE incorporates task-specific rewards that jointly enhance locatability assessment, visual clue reasoning, and geolocation accuracy. Both qualitative and quantitative results demonstrate that GLOBE outperforms state-of-the-art open-source LVLMs on geo-localization tasks, particularly in diverse visual scenes, while also generating more insightful and interpretable reasoning trajectories.
1 Introduction
The Background of Geo-localization. The rapid growth of visual content on social media and mobile devices, has made image geo-localization (determining where an image was taken) increasingly important for downstream applications such as autonomous navigation desouza2002vision and crisis response firmansyah2024empowering . Given that metadata (i.e., GPS coordinates) is frequently unavailable in practice friedland2010cybercasing , predicting geographic location from visual content remains a crucial capability. This demand has led to growing interest in the image geo-localization task hays2008im2gps .
Limitations in Existing Geo-localization Approaches. Traditional image geo-localization approaches fall into two main categories: classification and retrieval. Classification-based methods weyand2016 ; seo2018 ; muller2018 ; pramanick2022 ; clark2023 treat geo-localization as a discrete prediction task, assigning each image to a predefined set of geographical regions or cells. Retrieval-based methods yang2021cross ; zhu2022transgeo ; wang2023fine ; vivanco2023geoclip ; haas2024pigeon ; xia2025fg ; haas2023learning estimate location by comparing the query image to a large geo-tagged reference database, retrieving the closest match in terms of visual features, geographic coordinates, or semantic labels (e.g., city or country names). Although these methods perform well on standard benchmarks, they typically require training on millions of samples and lack interpretability, offering little insight into their underlying reasoning process.
When LVLMs Meet Geo-localization. The emergence of Large Vision-Language Models (LVLMs) liu2023llava ; Qwenvl ; Qwen2vl ; Qwen2.5vl ; chen2024internvl ; achiam2023gpt ; team2023gemini has introduced a new paradigm to tackle image geo-localization. Equipped with powerful multimodal reasoning capabilities and extensive world knowledge encoded through large-scale pretraining, LVLM-based methods li2024georeasoner ; jia2024g3 ; zhou2024img2loc have been explored through various strategies, including few-shot prompting, retrieval-augmented generation (RAG), and supervised fine-tuning (SFT). These methods are capable of generating both location predictions and explanations, offering greater interpretability in how decisions are made.
Figure 1: Overview of data and modeling limitations in LVLM-based image geo-localization.
Limitations in LVLM-based Image Geo-localization. Notably, geo-localization requires deeper reasoning than typical vision-language tasks. Success depends on more than recognition, as models must often draw on domain knowledge to infer plausible locations from subtle visual clues such as vegetation, architecture, or language, especially when iconic landmarks are absent. While LVLMs offer a promising path toward such reasoning-driven geo-localization, two fundamental challenges persist, as illustrated in Figure 1. On the data side, existing datasets rarely provide explicit reasoning supervision, such as interpretations of visual evidence and contextual justifications supporting the final location decision. Recent efforts li2024georeasoner ; dou2024gaga ; song2025geolocation to incorporate reasoning into geo-localization datasets have primarily relied on street-view imagery, which offers limited scene diversity and fixed viewpoints. As a result, models trained on such data often struggle to generalize to diverse, real-world visual conditions. On the modeling side, most current approaches depend on supervised fine-tuning with instruction-style data, which tends to encourage pattern replication rather than the development of a grounded understanding of visual-geographic relationships. Without verification mechanisms, these models rely heavily on correlation rather than structured inference, reducing their ability to generalize beyond familiar examples.
Figure 2: Example reasoning trajectories generated by GLOBE, illustrating interpretable and visually grounded geolocation predictions.
How GLOBE Tackles the Challenges. To address these challenges, we propose a novel pipeline for reasoning-aware geo-localization consisting of two main components: (1) constructing a geo-localization dataset from diverse social media images augmented with model-derived reasoning traces, and (2) fine-tuning a vision-language model using Group Relative Policy Optimization (GRPO) for enhanced reasoning. We begin by building MP16-Reason, an extension of MP-16 larson2017benchmarking , which contains user-captured photographs with diverse viewpoints and rich contextual content. To introduce reasoning supervision, we prompt multiple vision-language models Qwen2.5vl ; zhu2025internvl3 ; vivanco2023geoclip to distill the geolocation-related knowledge, including locatability assessments, reasoning trajectories, and predicted locations. To ensure the reliability of these distilled signals, we employ a multi-dimensional verification process that assesses both the alignment between visual evidence and model-generated reasoning, and the consistency across different models through self-verification, thereby filtering out inconsistent or hallucinated outputs. Finally, we fine-tune a pretrained LVLM on the curated dataset using GRPO shao2024deepseekmath , guided by task-specific rewards for locatability, visual grounding, and geolocation accuracy. Our resulting model, GLOBE, achieves state-of-the-art performance among open-source VLMs on geo-localization benchmarks, while producing more interpretable and visually grounded reasoning trajectories, as shown in Figure 2. Our main contributions include:
  • $\bullet$
    Reasoning-Aware Geo-Localization Dataset: We construct MP16-Reason, a diverse geo-localization dataset enriched with image-grounded reasoning supervision that supports model interpretability and generalization.
  • $\bullet$
    GRPO-Based Fine-Tuning: We develop a GRPO-based reinforcement learning framework that fine-tunes LVLMs using task-specific rewards for locatability, visual grounding, and geolocation accuracy, enabling stronger reasoning capabilities compared to traditional supervised fine-tuning.
  • $\bullet$
    Opensource LVLM: Trained through this pipeline, we opensource GLOBE. Empirical results demonstrate that GLOBE outperforms state-of-the-art LVLMs on multiple geo-localization benchmarks, while producing more interpretable and visually grounded reasoning trajectories.
2 Related Work
Image Geo-localization. Image geo-localization aims to predict the geographic location of a given image and has broad applications in urban analysis yeh1999urban ; firmansyah2024empowering ; yan2024urbanclip ; ye2019measuring ; ye2019visual , navigation desouza2002vision , and geospatial data mining korting2013geodma ; liu2024unitime ; liang2018geoman ; pan2019urban ; hao2025nature . With advances in multimodal models, research has evolved from classification weyand2016 ; seo2018 ; muller2018 ; pramanick2022 ; clark2023 and retrieval-based methods yang2021cross ; zhu2022transgeo ; wang2023fine ; vivanco2023geoclip ; haas2024pigeon ; xia2025fg ; haas2023learning to generation-based approaches li2024georeasoner ; zhou2024img2loc ; jia2024g3 , which aim to produce location predictions through visual reasoning. Recent studies li2024georeasoner ; zhou2024img2loc ; jia2024g3 have pointed out key limitations of classification (e.g., coarse granularity) and retrieval methods (e.g., dependency on large reference databases), prompting increased interest in generation-based alternatives. Since the introduction of the MediaEval Placing Tasks 2016 (MP-16) dataset by larson2017benchmarking , recent research zhou2024img2loc ; jia2024g3 continues to utilize this dataset to model relationships between visual semantics and geographic locations. In contrast to conventional approaches, current LVLMs liu2023llava ; Qwenvl ; Qwen2vl ; Qwen2.5vl ; chen2024internvl , which are typically pre-trained on large-scale datasets, inherently exhibit significant visual reasoning capabilities. This raises the critical question of whether the continued reliance on millions of labeled samples for supervised fine-tuning remains necessary to effectively adapt these models to specific tasks. In this work, we take a data-centric perspective to investigate how existing large-scale datasets can be leveraged to construct higher-quality training data for fine-tuning LVLMs in the context of image geo-localization.
Large Vision-Language Models. Building upon advancements in LLMs brown2020language ; touvron2023llama ; vicuna2023 ; qwen2023 ; bao2024harnessing ; guo2025deepseek ; he2025skywork , LLaVA liu2023llava was a pioneering approach in this direction, combining a pre-trained vision encoder with an LLM. It demonstrated that joint fine-tuning of visual and textual representations significantly enhances performance in tasks such as image-based question answering lu2022learn ; yue2024mmmu ; lumathvista ; li2024seed . Subsequently, various LVLMs have emerged Qwenvl ; Qwen2vl ; Qwen2.5vl ; chen2024internvl ; achiam2023gpt ; team2023gemini , differing primarily in their visual-language alignment mechanisms and associated architectural trade-offs. Motivated by these recent advancements, our work further investigates the shift of image geo-localization from traditional methods to LVLMs. Specifically, we explore how curated datasets can be effectively leveraged to facilitate more efficient fine-tuning of these models for geo-localization tasks.
Visual Reasoning and Verification. The emergence of advanced models such as DeepSeek liu2024deepseek has heightened expectations for the multimodal reasoning capabilities of LLMs. Most reasoning research imani2023mathprompter ; setlur2024rl has focused on mathematical tasks, with limited attention to open-ended or visual scenarios. Thus, these models often suffer from hallucination huang2025survey ; bai2024hallucination ; wei2024measuring , especially in visual tasks where they produce seemingly plausible but incorrect outputs. To address hallucination and promote more faithful reasoning, recent work has explored verification-based strategies lightman2023let ; dhuliawala2023chain ; lin2024fvel ; sriramanan2024llm ; oh2024erbench , as well as reinforcement learning frameworks shao2024deepseekmath ; zhou2025reinforced that optimize models via structured rewards. Motivated by these insights, we adopt GRPO as the reinforcement learning framework in our reasoning-driven geo-localization task.
3 [Uncaptioned image] GLOBE: The Methodology
We propose a novel pipeline based on the original MP-16 larson2017benchmarking dataset, aiming to advance image geo-localization from single-modal visual recognition to more robust multimodal reasoning. Achieving this objective requires not only powerful models but also well-curated training data that effectively capture geographic clues. Our pipeline for reasoning-aware geo-localization consists of two main components: dataset curation and model fine-tuning. These are implemented in three stages: (1) dataset curation via strong-to-weak distillation & verfication (Section 3.1), (2) reward construction via task-specific supervision (Section 3.2), and (3) model fine-tuning via GRPO-based reinforcement learning (Section 3.3).
3.1 Dataset Curation: Data Distillation & Verfication
Raw web-scale datasets contain a diverse range of social media images captured from varied perspectives. However, these datasets suffer from substantial noise deng2009imagenet ; li2017webvision ; wei2022learning ; wei2023aggregate ; liu2024automatic , such as close-up shots with limited visual context or generic objects lacking informative localizable clues. To address this issue and select appropriate images for downstream training, we employ multiple vision-language model knowledge distillation for data synthesis and multi-dimensional verification for data curation.
Figure 3: The pipeline of data synthesis and curation via multi-model distillation and verification.
Multiple Vision-Language Models Knowledge Distillation. We utilize multiple vision-language models (e.g., Qwen2.5-VL-72B Qwen2.5vl , InternVL3-78B zhu2025internvl3 , and GeoCLIP vivanco2023geoclip ) to extract locatability judgments, visual clues, and geolocation predictions for each image in the MP-16 larson2017benchmarking dataset, following the approach of li2024georeasoner ; campos2025gaea . As shown in Figure 3, Qwen2.5-VL and InternVL3 produce binary locatability decisions, step-by-step reasoning trajectories, and textual geolocation predictions. GeoCLIP, in contrast, produces latitude-longitude coordinates along with a confidence score that quantifies locatability campos2025gaea . Collectively, these strong models offer complementary signals, which we distill into structured supervision for downstream data curation and reward modeling.
Multi-dimensional Verfication. Following model inference, we perform multi-dimensional verification to curate high-quality data, as illustrated in Figure 3. Initially, we filter out images with negative locatability decisions or low locatability scores. Subsequently, incorrect geolocation predictions are discarded by comparing them against ground-truth annotations. To ensure the reliability of the knowledge distilled from Qwen2.5-VL and InternVL3, we introduce a self-verification step in which the geolocation predictions and reasoning trajectories of both models are compared for each image. Only those samples exhibiting consistent location outputs (e.g., matching city- or country-level predictions) and semantically aligned reasoning chains are retained. This cross-model agreement serves as the reliability proxy in distilled supervision. Furthermore, to enforce visual grounding of the reasoning process, we employ a general-purpose semantic segmentation model cheng2021per to extract both the categories and relative proportions of visual elements within each image. We then assess the consistency between the entities mentioned in the reasoning trajectories and the visual elements identified via segmentation. Through this multi-stage validation pipeline, which combines locatability filtering, self-verification of distilled knowledge, and visual-semantic consistency checks, we curate a robust and trustworthy dataset tailored for downstream tasks.
3.2 Reward Construction: Task-specific Supervision
Building upon the curated dataset introduced in Section 3.1, we develop three task-specific rewards to assess distinct dimensions of reasoning quality in the geo-localization process. Each reward is trained with annotated supervision and collectively provides a structured reward signal, which guides the policy optimization during the reinforcement learning stage described in Section 3.3.
Formally, let $\mathcal{D}={(I_{i},y_{i},g_{i},r_{i})}_{i=1}^{N}$ denote the curated dataset of $N$ samples, where $I_{i}$ is an image, $y_{i}\in\{0,1\}$ is a binary label indicating whether the image is localizable, $g_{i}$ indicates the ground-truth geolocation, and $r_{i}$ is the associated reasoning trajectory.
Locatability Reward. We develop a binary classification reward model to estimate the localizability of an image based on its visual content. Using the curated dataset $\mathcal{D}$, we train a model $R_{\text{loc}}(I_{i})\in\{0,1\}$ to predict the probability that $y_{i}=1$, i.e., the image is localizable. Accordingly, we define the reward as:
$R_{\text{loc}}(I_{i})=\mathbb{P}(y_{i}=1\mid I_{i};\theta_{\text{loc}}),$ (1)
where $\theta_{\text{loc}}$ denotes the parameters of the reward model. The output score serves both as a reward signal for reinforcement learning and as a soft indicator of an image’s inherent locatability.
Visual Grounding Consistency Reward. To ensure the model-generated reasoning aligns with the actual visual content, we introduce a reward model evaluating entity grounding consistency. For a given sample $(I_{i},r_{i})$ from the curated dataset, let $\hat{r}_{i}$ denote the predicted reasoning. We extract a set of entities $E_{i}=\{e_{1},e_{2},...,e_{n}\}$ from the reasoning trajectory $\hat{r}_{i}$, and a set of visual elements $V_{i}=\{v_{1},v_{2},...,v_{m}\}$ from both the image $I_{i}$ (via semantic segmentation) and the text of $r_{i}$ (via entity extraction). We define a soft matching function $\text{Match}(e_{j},V_{i})\in{0,1}$, which returns 1 if entity $e_{j}$ approximately matches any element in $V_{i}$, allowing for partial lexical or semantic overlap. The visual grounding reward is computed as:
$R_{\text{vis}}(I_{i},\hat{r}_{i},r_{i})=\frac{1}{|E_{i}|}\sum_{j=1}^{|E_{i}|}\text{Match}(e_{j},V_{i}),$ (2)
where $R_{\text{vis}}$ assigns a higher score when more entities in the reasoning are visually grounded. This reward penalizes hallucinated entities that do not correspond to visible elements in the image, thereby encouraging grounded visual reasoning.
Geo-localization Accuracy Reward. To evaluate model predictions at a semantic location level, we define a classification-based reward that reflects whether the predicted country and city match the ground truth. Let $\hat{g}_{i}=(\hat{c}_{i},\hat{t}_{i})$ denote the predicted country and city for image $I_{i}$, and let $g_{i}=(c_{i},t_{i})$ be the corresponding ground-truth geolocation from the curated dataset. The geo-localization reward $R_{\text{geo}}$ is defined as:
$R_{\text{geo}}(\hat{g}_{i},g_{i})=\mathbb{I}[\hat{c}_{i}=c_{i}]\cdot\left(\alpha\cdot\mathbb{I}[\hat{t}_{i}=t_{i}]+(1-\alpha)\right),$ (3)
where $\mathbb{I}[\cdot]$ is the indicator function and $\alpha\in[0,1]$ is a weighting factor that controls the importance of city-level correctness, conditional on the country being correct. This reward structure captures the hierarchical nature of geo-tags. A reward of 00 is assigned when the predicted country is incorrect (i.e., $\hat{c}_{i}\neq c_{i}$). If the country is correct but the city is not (i.e., $\hat{c}_{i}=c_{i}$, $\hat{t}_{i}\neq t_{i}$), the model receives a partial reward of $1-\alpha$. A full reward of $1$ is assigned only when both predictions are correct (i.e., $\hat{c}_{i}=c_{i}$, $\hat{t}_{i}=t_{i}$). This tiered design encourages the model to first learn coarse-grained localization before refining its predictions to finer spatial resolutions.
Figure 4: GRPO optimization framework with multi-dimensional reward design. For each prompt, candidate outputs are scored using three task-specific reward models: $R_{\text{loc}}$, $R_{\text{vis}}$, and $R_{\text{geo}}$, which reflect different aspects of geo-localization reasoning. Group-wise advantage values guide policy updates, while a $\mathcal{D}_{\text{KL}}$ penalty constrains divergence from the reference model.
3.3 Model Fine-tuning: GRPO-based Reinforcement Learning
With the reward signals defined in Section 3.2, we fine-tune the base model using GRPO shao2024deepseekmath , a reinforcement learning algorithm designed for ranking-based reward optimization, as illustrated in Figure 4. GRPO builds upon Proximal Policy Optimization (PPO)schulman2017proximal , which stabilizes policy updates by optimizing a clipped surrogate objective using advantage estimates derived from scalar rewards. Unlike PPO, GRPO introduces group-wise normalization and optimizes relative preferences among candidates conditioned on each prompt, enhancing robustness to variations in the reward scale.
Let $\pi_{\theta}$ denote the current policy parameterized by $\theta$, and let =$\mathcal{B}=\{(\bm{x}_{i},\{\bm{a}_{i}^{(j)}\}_{j=1}^{k})\}$ represent a batch of input prompts $\bm{x}_{i}$ each paired with $k$ candidate completions $\bm{a}_{i}^{(j)}$ sampled from the policy. Each completion $\bm{a}_{i}^{(j)}$ is scored by a composite reward function:
$r_{i}^{(j)}=\lambda_{1}R_{\text{loc}}+\lambda_{2}R_{\text{vis}}+\lambda_{3}R_{\text{geo}},$ (4)
where $\lambda_{1},\lambda_{2},\lambda_{3}\in[0,1]$ are weights controlling the importance of the three reward components: locatability ($R_{\text{loc}}$), visual grounding consistency ($R_{\text{vis}}$), and geo-localization accuracy ($R_{\text{geo}}$).
To encourage the model to prefer higher-reward completions within each group, GRPO computes a group-normalized advantage for each candidate:
$A_{i}^{(j)}=r_{i}^{(j)}-\frac{1}{k}\sum_{l=1}^{k}r_{i}^{(l)},$ (5)
which centers rewards within each prompt group. Eqn. (5) guides the policy to optimize relative ranking rather than absolute scores, making it suitable for scenarios with non-uniform reward scales.
The policy is then updated by maximizing the following clipped surrogate objective:
$\mathcal{L}_{\text{GRPO}}(\theta)=\mathbb{E}_{(\bm{x}_{i},\bm{a}_{i}^{(j)})\sim\pi_{\theta_{\text{ref}}}}\left[\min\left(\rho_{i}^{(j)}A_{i}^{(j)},\text{clip}(\rho_{i}^{(j)},1-\epsilon,1+\epsilon)A_{i}^{(j)}\right)-\beta\mathcal{D}_{\text{KL}}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]\right],$ (6)
where $\rho_{i}^{(j)}=\frac{\pi_{\theta}(\bm{a}_{i}^{(j)}|\bm{x}_{i})}{\pi_{\theta_{\text{old}}}(\bm{a}_{i}^{(j)}|\bm{x}_{i})}$ is the likelihood ratio between the current and reference policies, and $\epsilon$ is the clipping threshold. The coefficient $\beta$ controls the strength of the $\mathcal{D}_{\text{KL}}$ penalty, and $\pi_{\text{ref}}$ is the reference policy used to constrain updates. In practice, the reference policy $\pi_{\text{ref}}$ is typically instantiated as the previous policy snapshot, serving to regularize updates and ensure training stability.
4 Experiments
We conduct qualitative and quantitative experiments to evaluate the effectiveness of our curated dataset MP16-Reason and the GRPO-based training strategy used in GLOBE. Specifically, we examine whether MP16-Reason enables better geo-reasoning (i.e., the ability to infer geographic locations through interpretable and visually grounded reasoning) compared to conventional image-only datasets (which lack reasoning supervision) and street-view datasets (which offer limited visual diversity), and whether GRPO training yields stronger reasoning performance than supervised fine-tuning. We also compare GLOBE with both open- and closed-source LVLMs.
4.1 Experimental Setup
Datasets. The curated dataset MP16-Reason is divided into two subsets: MP16-Reason-Train with 33k samples and MP16-Reason-Test with 12k samples, respectively. MP16-Reason-Train is used to train GLOBE, while MP16-Reason-Test is used to evaluate all baseline methods. To ensure a comprehensive comparison, we additionally evaluate all models on the public geo-localization benchmark IM2GPS3K vo2017revisiting .
Evaluation Metrics. We adopt different evaluation metrics for MP16-Reason-Test and the public geo-localization benchmark to account for differences in annotation format and prediction targets. In MP16-Reason-Test, we evaluate model performance using three metrics: (1) city-level accuracy, (2) country-level accuracy. On the public benchmark, we follow previous work vivanco2023geoclip ; haas2024pigeon ; jia2024g3 ; li2024georeasoner and report the percentage of predictions whose geographic distance to the ground-truth coordinate falls within fixed thresholds (25km, 200km, and 750km). Since our model outputs discrete place names (e.g., country or city), we use external tools to convert these names into their corresponding geographic center coordinates for evaluation.
Implementation details. We implement GLOBE based on the publicly available Qwen2.5-VL-7B Qwen2.5vl , a large vision-language model with strong multimodal understanding capabilities. Instead of using task-specific supervised fine-tuning as a cold start, we directly fine-tune the model using reinforcement learning based on the GRPO framework described in Section 3.3. All training is conducted using MP16-Reason-Train. Further details are provided in Appendix A.1.
Table 1: Localization accuracy on MP16-Reason-Test and IM2GPS3K vo2017revisiting . Underlined results indicate test–train overlap; † denotes models that are not publicly available. Best and second-best results are in bold and blue, respectively.
Method Dataset Size MP16-Reason Test IM2GPS3K vo2017revisiting  (@ km)
City Country City Region Country
Acc. Acc. 25km 200km 750km
ISNs muller2018 MP-16 4M 47.33 82.06 28.00 36.60 49.70
GeoCLIP vivanco2023geoclip MP-16 4M 52.52 84.07 34.47 50.65 69.67
Translocator pramanick2022 MP-16 4M - - 31.10 46.70 58.90
PIGEOTTO haas2024pigeon MP-16 4M - - 36.70 53.80 72.40
G3 (LLaVA) jia2024g3 MP-16 4M - - 35.87 49.42 66.93
G3 (GPT4V) jia2024g3 MP-16 4M - - 40.94 55.56 71.24
Qwen2.5-VL-7B Qwen2.5vl - - 37.33 78.38 32.53 43.14 58.96
Gemma3-27B team2025gemma - - 37.12 72.47 33.37 46.61 63.63
InternVL3-78B zhu2025internvl3 - - 37.31 78.86 35.05 47.32 64.03
Qwen2.5-VL-72B Qwen2.5vl - - 44.45 82.97 35.77 48.35 64.96
GPT-4.1 gpt4-1 - - 53.78 89.87 46.85 60.36 74.41
GeoReasoner li2024georeasoner GSV 133K 32.39 67.30 26.94 36.63 52.27
GaGA dou2024gaga MG-Geo 5M - - 33.00 48.00 67.10
GLOBE (Ours) MP16-Reason 33K 53.16 88.54 40.18 56.19 71.45
4.2 Performance Comparison
We evaluate the effects of the curated dataset and training strategy through quantitative metrics and qualitative analysis, as detailed in the following subsections.
4.2.1 MP16-Reason: Annotation Benefit and Scene Robustness
We validate the contribution of the MP16-Reason dataset through two perspectives. First, we examine whether reasoning-augmented annotations provide tangible benefits over image-only supervision. Second, we compare MP16-Reason to existing geo-reasoning datasets that are predominantly street-view-centric, assessing generalization performance across diverse visual scenes.
Annotation Benefit. To isolate the contribution of our reasoning-based annotations, we ablate the reward components used in GRPO training. Specifically, we compare a model trained only with the geo-localization accuracy reward (VGC Reward), which is derived from the original annotations, as shown in Table 2-row4. As reported in Table 1 and Table 2, GLOBE trained on MP16-Reason significantly outperforms all variants trained on MP-16 (Table 1-row1, Table 1-row2, and Table 1-row3). While PIGEOTTO haas2024pigeon outperforms GLOBE at the country level, its performance lags significantly at the city and region levels (Table 1-row4). It is worth noting that G3 (GPT-4V) jia2024g3 leverages GPT-4V to perform retrieval-augmented generation (RAG) for geo-localization, incorporating external knowledge through a prompt ensembling strategy. Despite this advantage, our method achieves comparable performance at the city level and even slightly outperforms G3 (GPT-4V) at the region and country levels, as shown in Table 1. This highlights the importance of the additional reasoning-based annotations, such as locatability judgments and visual grounding signals, which were distilled from multiple vision-language models and used to supervise finer-grained aspects of geo-reasoning.
Scene Robustness. To assess the generalization ability of geo-reasoning models beyond street-view scenarios, we compare GLOBE with counterpart LVLMs trained on street-view-oriented reasoning datasets (Table 1-row12 and Table 1-row13). While models trained on such datasets perform well on street-view imagery, their generalization degrades considerably in more diverse scenarios (see Table 1). In contrast, MP16-Reason, which includes both street-view and non-street-view images, achieves consistently strong results across a wider range of visual contexts, better reflecting the complexity of real-world geo-localization tasks.
Figure 5: Reasoning comparison of four different models (GPT-4.1 gpt4-1 , GLOBE, Qwen2.5-VL-7B Qwen2.5vl with SFT, and InternVL3-78B zhu2025internvl3 ) on the same input image. Reliable visual clues identified by the models are marked in text.
4.2.2 [Uncaptioned image] GLOBE: Model Performance and Training Strategy Analysis
To further evaluate the effectiveness of our proposed model GLOBE, we conduct comprehensive experiments from two perspectives: external baseline comparison and internal strategy ablation. Externally, we benchmark GLOBE against vision-only models without reasoning supervision, general-purpose LVLMs, and task-specific LVLMs trained on street-view reasoning datasets. Internally, we benchmark SFT against GRPO-based reinforcement learning, and perform ablation studies within the GRPO framework to assess the effect of different reward configurations. These comparisons examine the performance of GLOBE and the effects of its key design choices.
External Baseline Comparison. We compare GLOBE against three categories of representative baselines. The first group includes traditional approaches, which rely on visual feature matching without supporting reasoning capabilities. The second group consists of open-source LVLMs (e.g., Qwen2.5-VL Qwen2.5vl , InternVL3 zhu2025internvl3 , Gemma3 team2025gemma , and GPT-4.1 gpt4-1 ) trained on general-purpose data, which are expected to exhibit broader and more generalizable reasoning capabilities. The third group includes task-specific LVLMs trained on geo-reasoning datasets oriented toward street-view imagery. As shown in Table 1, GPT-4.1 gpt4-1 demonstrates outstanding performance on this task, but it remains closed-source. Aside from GPT-4.1 gpt4-1 , the proposed GLOBE achieves strong accuracy on both the MP16-Reason-Test and the public benchmark IM2GPS3K vo2017revisiting , while also producing more coherent and interpretable reasoning trajectories (see Table 1 and Figure 5). Notably, GLOBE achieves this performance with just 33K samples from MP16-Reason, highlighting the efficiency of reasoning-aware supervision.
Table 2: Ablation study on training methods and reward modeling configurations in MP16-Reason-Test set using Qwen2.5-VL-7B-Instruct model.
Models CoT SFT GRPO Metrics
Loc Reward VGC Reward GA Reward City Country
Acc. Acc.
Qwen2.5-VL-7B 37.21 78.28
Qwen2.5-VL-7B 46.38 84.92
GLOBE w/o Loc&GA 48.80 83.65
GLOBE w/o Loc&VGC 50.27 87.31
GLOBE w/o Loc 50.47 87.15
GLOBE w/o VGC 51.68 87.62
GLOBE w/o GA 49.11 86.53
GLOBE 53.16 88.54
Internal Strategy Ablation. We further investigate the impact of different training paradigms and reward configurations on the performance of GLOBE. Specifically, we compare full-parameter SFT, and our proposed GRPO-based reinforcement learning approach (See Table 2-row2 and Table 2-row8). For GRPO, we conduct a reward ablation by evaluating different combinations of the three reward components: Locatability (Loc) Reward, Visual Grounding Consistency (VGC) Reward, and Geo-localization Accuracy (GA) Reward. It demonstrates that GRPO with the complete reward set achieves the highest overall performance (Table 2-row8). Removing either Loc, VGC or GA Reward results in noticeable drops, underscoring the importance of reasoning-aware supervision beyond location-level correctness (Table 2-row5, Table 2-row6, Table 2-row7). In addition, GRPO demonstrates clear advantages over SFT for the image geo-localization task in LVLMs, offering improved consistency and grounding by directly optimizing the relative quality of generated outputs (Table 2-row2 vs. Table 2-row8).
5 Discussion
Toward Fine-Grained Geo-localization: Limits of Pure Reasoning. While our reasoning-aware framework achieves strong performance at the country and city levels, its effectiveness diminishes when tasked with fine-grained, coordinate-level localization. This limitation originates from the inherent nature of the reasoning process: predictions are based on high-level semantic cues such as language, architectural style, or vegetation, which often lack the spatial specificity required to differentiate between closely situated locations. For example, multiple European cities may share similar visual patterns, such as Mediterranean-style architecture, the presence of European Union flags, or public signage in English, which makes it difficult for the model to resolve fine-grained geographic ambiguities through reasoning alone. In such cases, even accurate reasoning can only narrow down a broad region but cannot pinpoint an exact location. This highlights a key challenge in reasoning-centric geo-localization: the lack of precise visual-geographic anchoring. To overcome this limitation, future work may explore hybrid approaches that combine reasoning to constrain the candidate region, followed by local feature-based retrieval within that region to achieve coordinate-level precision.
Beyond Scale Alone: Data Efficiency in Reasoning-aware Training. Our experiments show that training GLOBE on just 33K high-quality, reasoning-aware samples (MP16-Reason) achieves performance comparable to, and sometimes exceeding, models trained on millions of generic image-text pairs. This highlights that for reasoning-centric tasks, targeted supervision can be more effective than sheer data scale. Our results suggest that aligning supervision with task-specific reasoning offers a more data-efficient path forward for LVLM training.
Beyond Geo-localization: GRPO for Reasoning-driven LVLM Tasks. Our findings suggest that GRPO, as a training paradigm, is particularly well-suited for reasoning-driven objectives in LVLMs. Unlike SFT, which often treats outputs as isolated targets, GRPO directly optimizes the relative quality of outputs through scalar reward signals. This form of supervision allows GRPO to guide complex reasoning behaviors in a more structured and interpretable manner than traditional training objectives. While our work focuses on geo-localization, we believe the GRPO paradigm can be readily extended to other multimodal reasoning tasks, such as visual question answering and multimodal chain-of-thought generation.
6 Conclusion
In this paper, we present a novel reasoning-driven pipeline for image geo-localization by leveraging LVLMs. To address the limitations of existing datasets and training paradigms, we introduce MP16-Reason, a high-quality dataset constructed from diverse social media images and enriched with automatically distilled locatability labels and reasoning trajectories. Building upon this dataset, we propose GLOBE, an LVLM trained via GRPO-based reinforcement learning, which jointly improves three core aspects of geo-localization: locatability assessment, visual clue reasoning, and geo-location recognition. In contrast to SFT, our GRPO-based training framework directly optimizes reasoning quality through structured reward signals, leading to substantial gains in both interpretability and localization accuracy. Empirical results show that GLOBE, using only 33K data, achieves performance comparable to or better than state-of-the-art methods trained on millions of samples.
Appendix A Appendix
A.1 Implementation Details
A.1.1 Prompt Template
To ensure fairness and consistency in both training and evaluation, we standardize the input format for all LVLMs. Specifically, images are fed into the model along with a unified prompt template. This prompt is designed to guide the model toward reasoning about geographic clues in a consistent manner across different stages and model variants. During both fine-tuning and testing, we use the same prompt format to minimize prompt-induced variance and allow for fair comparison of model performance.
Prompt Template You are a geolocation expert. You are participating in a geolocation challenge. Based on the provided image: 1. Carefully analyze the image for clues about its location (architecture, signage, vegetation, terrain, etc.) 2. Think step-by-step about what country, and city this is likely to be in and why Your final answer include these two lines somewhere in your response: country: [country name] city: [city name] You MUST output the thinking process in <think> </think> and give answer in <answer> </answer> tags.
A.1.2 Training Data Example
The example below illustrates a training sample from MP16-Reason. Each instance contains the image path, a reasoning trajectory generated by vision-language models, ground-truth city and country labels, and a set of grounded entities extracted from the reasoning text.
Training Data Example img_path: "3f_e4_302010632.jpg" reason: "The image shows a large stadium filled with spectators wearing red clothing, indicating a strong team spirit and possibly a sports event. There is a scoreboard displaying ‘Ohio State’ and ‘Michigan’, which are well-known college football teams from the United States. This suggests that the event is an American college football game. The presence of cheerleaders, a marching band, and the large crowd further supports the idea that this is a major sporting event in the United States. The architecture of the stadium, including the design of the scoreboard and the layout of the seating, is consistent with large American college football stadiums. Given the reference to ‘Ohio State’, it is highly likely that this event is taking place in Ohio, specifically at Ohio State University’s home stadium, which is located in Columbus." city: "Columbus" country: "United States" entities: [{"text": "stadium", "type": "ARCH"}, {"text": "scoreboard", "type": "SIGN"}, {"text": "seating", "type": "ARCH"}]
A.1.3 Hyper-parameter Settings
We summarize the key hyper-parameters used in training GLOBE in Table 3. These settings are selected based on standard practices in fine-tuning large vision-language models and further adjusted through preliminary ablation studies on a held-out validation set. Unless otherwise specified, all experiments are conducted using the same configuration to ensure comparability and reproducibility.
Table 3: The hyper-parameter settings of the proposed GLOBE.
Hyper Params Value
Learning Rate 1e-6
Total Batch Size 16
Weight Decay 0.1
Warmup Ratio 0.01
Optimizer AdamW
Adam Beta1 0.9
Adam Beta2 0.95
LR Scheduler cosine
Model Max Length 8192
A.2 Limitation
While our reasoning-based framework performs well at country and city levels, its accuracy declines in fine-grained, coordinate-level geo-localization. This is due to the abstract nature of reasoning, which relies on high-level semantic cues (e.g., architecture, language, vegetation) that often lack the spatial precision needed to distinguish between visually similar, nearby locations. As a result, even correct reasoning may only localize to a broad region. Future work could address this by combining reasoning with local feature-based retrieval to improve fine-grained accuracy.
A.3 Broader Impacts
This work introduces a reasoning-oriented geo-localization framework that leverages diverse social media imagery and bi-objective optimization to enhance the reasoning capabilities of large vision-language models. While this approach improves interpretability and performance in complex visual scenes, it also raises privacy and misuse concerns. The ability to infer precise locations from user-shared images may lead to unauthorized tracking, surveillance, or profiling, especially if deployed at scale without appropriate safeguards.
To mitigate these risks, we recommend restricting access to the model via gated APIs, incorporating uncertainty estimation in predictions, and clearly documenting limitations and intended use cases. Special care should be taken when applying the method to user-generated content, including adherence to data licenses and privacy-preserving practices. Responsible deployment will be essential to ensure the benefits of improved geo-reasoning do not come at the cost of societal harm.